OpenStudy (anonymous):

The equation of a line is y=-1/2x-2. What is the equation of the line that is perpendicular to the first line and passes through the point (2, –2)?

OpenStudy (anonymous):

A. y = 2x – 2 B. y=1/2x-3 C. y = 2x – 6 D. y = –2x + 2

OpenStudy (anonymous):

y = 2x-6

OpenStudy (anonymous):

Thanks, but could you explain the process to me?